number of computers sold per day at a local computer store are 0,1,2,3, and 4 with the probabilities 0.1, 0.2, 0.3, 0.2, and 0.2 respectively. Find the variance and standard deviation of the distribution


1
Expert's answer
2022-02-28T17:55:12-0500

we first obtain the mean as below

mean= ( (0.1*0) + (0.2 *1) + (0.3 *2) + (0.2*3) + (0.2*4) ) = 2.2

Thus variance =( (0.1 *02) + (0.2 *12) + (0.3 * 22) + (0.2*32) +(0.2*42) ) - (mean)2

= 6.4 - 4.84

= 1.56

standard deviation = (variance)1/2 = (1.56)1/2 = 1.2489996